### v1

# Multimodal RewardBench 2: Evaluating Omni Reward Models for Interleaved Text and Image

Abstract
Reward models (RMs) are essential for training large language models (LLMs), but remain underexplored for omni models that handle interleaved image and text sequences. We introduce Multimodal RewardBench 2 (MMRB2), the first comprehensive benchmark for reward models on multimodal understanding and (interleaved) generation. MMRB2 spans four tasks: text-to-image, image editing, interleaved generation, and multimodal reasoning (“thinking-with-images”), providing 1,000 expert-annotated preference pairs per task from 23 models and agents across 21 source tasks. MMRB2 is designed with: (1) practical but challenging prompts; (2) responses from state-of-the-art models and agents; and (3) preference pairs with strong human-expert consensus, curated via an ensemble filtering strategy. Using MMRB2, we study existing judges for each subtask, including multimodal LLM-as-a-judge and models trained with human preferences. The latest Gemini 3 Pro attains 75-80% accuracy. GPT-5 and Gemini 2.5 Pro reach 66–75% accuracy, compared to >90% for humans, yet surpass the widely used GPT-4o (59%). The best performing open-source model Qwen3-VL-32B achieves similar accuracies as Gemini 2.5 Flash (64%). We also show that MMRB2 performance strongly correlates with downstream task success using Best-of-N sampling and conduct an in-depth analysis that shows key areas to improve the reward models going forward.
[…]
1 Introduction
Figure 1: Examples of multimodal preference pairs in MMRB2 across four subtasks: text-to-image generation, interleaved generation, image editing, and multimodal reasoning, showing human and model judgments on challenging prompts.
Reward models are central to the development of LLMs (Christiano et al., 2017; Bai et al., 2022; Jaech et al., 2024; Guo et al., 2025; Lambert et al., 2024; Yuan et al., 2024). They enable scalable evaluation that tracks model performance and surfaces systematic weaknesses (Zheng et al., 2023). They can be used to assess data quality, which is crucial for building synthetic data pipelines (Wang et al., 2022b). And, as reinforcement learning becomes increasingly important in post training, high quality reward models are crucial for surfacing or suppressing a range of different model capabilities (Christiano et al., 2017; Wu et al., 2023c; Guo et al., 2025). Recent work has focused on developing new classes of omni models, which enable understanding, generation, and reasoning with interleaved text and images (OpenAI, 2024; Chameleon Team, 2024; Ge et al., 2025; Zhou et al., 2024; Deng et al., 2025; Chen et al., 2025c; Wang et al., 2024b; Chen et al., 2025a; Google DeepMind, 2025a). However, reward modeling for omni models remains largely unexplored.
This omission is at least in part because there is no existing benchmark for omni reward models, making it nearly impossible to measure model quality. Unlike text-only models, omni models can generate and understand any number of texts and images together in a single arbitrarily ordered sequence. This generality creates unique challenges for reward modeling. Unlike domains such as math or coding, images are difficult to verify automatically (Hessel et al., 2021; Hu et al., 2023; Lin et al., 2024), and high-quality preference data requires carefully designed annotation protocols (Liang et al., 2024). Omni models can also be used for a very broad range of real-world applications, demanding diverse task coverage for both training and evaluation (Liu et al., 2024; Chen et al., 2024; Yao et al., 2025). Finally, gathering high-quality responses needed to train and evaluate omni reward models can be challenging, since omni model capabilities are not always as strong as the models used to develop previous text-only reward benches.
We introduce Multimodal RewardBench 2 (MMRB2) which overcomes all of these challenges to establish a foundation for future research on omni reward modeling. MMRB2 follows Multimodal RewardBench (MMRB1) (Yasunaga et al., 2025), which covered image-text-to-text tasks for multimodal large language models (MLLMs). MMRB2 instead covers the much more challenging case of omni models over four subtasks (Figure 1): text-to-image, image editing, interleaved generation, and multimodal reasoning (“thinking with images” (OpenAI, 2025d)). Each subtask contains 1,000 expert-annotated preference pairs, consisting of a task prompt, a preferred response, and a rejected response. To ensure that MMRB2 is comprehensive, reliable, and highly predictive of reward model quality, we design it with three key characteristics: (1) diverse, practical, yet challenging prompts near the capability boundary of frontier models, drawn from 21 existing and newly created tasks; (2) responses generated by state-of-the-art multimodal models, ranging from SD3.5 (Stability AI, 2024) to GPT-Image (OpenAI, 2025c) and Gemini 2.5 Flash Image (Google DeepMind, 2025a), along with specialized agents (Hu et al., 2024) for interleaved generation and visual reasoning tasks where even the best models often fail; and (3) preference pairs that have >90% agreement among human experts but which remain challenging for current judges (both MLLM-based judges and trained reward models), curated via an ensemble filtering strategy. A summary of all the prompts and multimodal models covered in MMRB2 are in Table 1.
Category Source Response Models Task Description Text-to-Image WISE (Niu et al., 2025)
EvalMuse (Han et al., 2024)
OneIG-Bench (Chang et al., 2025)
R2IBench (Chen et al., 2025b)
RealUnify (Shi et al., 2025)
Image Gemini 2.0 and 2.5 Flash Image (Google DeepMind, 2025c, a)
Imagen 3 (Baldridge et al., 2024)
Imagen 4 and Ultra (Google DeepMind, 2025d)
FLUX (Labs et al., 2025)
GPT-image-1 (OpenAI, 2025c)
SD 3.5-L (Stability AI, 2024)
Image generation from text assessing creativity, composition, reasoning, text rendering, etc.
Image Editing DreamBench (Peng et al., 2025)
Emu-Edit (Sheynin et al., 2024)
HQ-Edit (Hui et al., 2024)
RISE-Bench (Zhao et al., 2025)
Text-heavy edit
Multi-Image edit
Image Gemini-2.0 and 2.5 Flash Image (Google DeepMind, 2025c, a)
Imagen3-Edit (Baldridge et al., 2024)
FLUX-Kontext (Labs et al., 2025)
GPT-image-1 (OpenAI, 2025c)
Object replacement, scene modification, style change, entity-preserving editing, reasoning-heavy editing, text-heavy editing, multi-image editing, etc.
Interleaved Generation Chameleon (Chameleon Team, 2024)
Interleaved-Eval (Liu et al., 2024)
ISG-Bench (Chen et al., 2024)
MMMG (Yao et al., 2025)
Text+Image Gemini 2.0 and 2.5 Flash Image (Google DeepMind, 2025c, a)
Agents:
GPT-Gemini-agent
GPT-GPT-image-agent
GPT-Imagen-agent
GPT-FLUX-agent
Interleaved text-image generation, storytelling, open-ended visual question answering, scene composition, 3D prediction, temporal prediction, etc.
Reasoning BLINK (Fu et al., 2024)
MindCube (Yin et al., 2025)
VisuLogic (Xu et al., 2025)
V (Wu and Xie, 2023)
MuirBench (Wang et al., 2024a)
RealUnify (Shi et al., 2025)
Text(+Image) GPT-5 (OpenAI, 2025b)
GPT-4.1 (OpenAI, 2025a)
GPT-4o (OpenAI, 2024)
o3 (OpenAI, 2025d)
Gemini 2.5 Flash (Gemini Team, 2025)
Gemini 2.5 Pro (Gemini Team, 2025)
Sketchpad Agents (Hu et al., 2024):
o3-sketchpad-agent
GPT-5-sketchpad-agent
Thinking with images, spatial reasoning, multi-image reasoning, perception-heavy tasks, etc.
Table 1: Overview of the four subtask categories in MMRB2, including their prompt sources, response modalities, model that were used to synthesize the data, and task descriptions. The benchmark draws from a diverse set of public and newly created datasets to cover text-to-image generation, image editing, interleaved text–image generation, and multimodal reasoning ("thinking with images").
Using MMRB2, we conduct a comprehensive study of reward models for multimodal understanding and generation, including multimodal LLM-as-a-judge, task-specific metrics, and reward models trained with human preferences. Experiments show that:
  • MMRB2 poses significant challenges to current MLLM-as-a-judge approaches: the latest Gemini 3 Pro (Google DeepMind, 2025b) reaches 74-80% accuracy across all subtasks. GPT-5 (OpenAI, 2025b) and Gemini 2.5 Pro (Gemini Team, 2025) achieve only moderate performance (66-75% accuracy across all subtasks) compared to >90% for humans. The best open-source model, Qwen3-VL-32B (Qwen Team, 2025), achieves 55%-69% accuracy. Notably, GPT-4o (OpenAI, 2024), which is commonly used as an evaluator in existing benchmarks, attains only 51-65% accuracy, suggesting that it is no longer suitable for evaluating frontier multimodal models, especially on reasoning-heavy tasks.
  • We study task-specific metrics (e.g., VQAScore (Lin et al., 2024)) and reward models trained on human preferences (e.g., ImageReward (Xu et al., 2023), UnifiedReward (Wang et al., 2025)), and find that they are no longer reliable on the challenging prompts and frontier models in MMRB2. For instance, VQAScore (with Qwen2.5-VL backbone) and ImageReward achieve 58.3% and 54.0% on text-to-image evaluation, respectively, well below MLLM-as-a-judge baselines such as Qwen3-VL-32B (64.1%) and Gemini 3 Pro (74.4%). While human preference training improves performance over heuristic metrics, these models still fall short of frontier MLLMs.
  • We show that performance on MMRB2 strongly correlates with performance on GenAI-Bench (Li et al., 2024), GEdit-Bench (Liu et al., 2025), ISGBench (Chen et al., 2024), and EMMA (Hao et al., 2025) when using different reward models for best-of-$N$ selection, suggesting that MMRB2 is a good proxy for downstream effectiveness.
  • Further analysis of test-time scaling and fine-grained error patterns reveals substantial remaining headroom for omni model reward modeling and highlights concrete failure modes that future methods should address. Judges show notably higher agreement with human preferences on different-model pairs than on same-model pairs, with differences of up to 12%. Moreover, in multimodal reasoning tasks, judges exhibit a strong bias toward responses that include images, with performance gaps of 27.7–49.3% between pairs where annotators preferred image-containing responses and those where the preferred response contained only text.
Overall, MMRB2 establishes a challenging and informative benchmark that we hope will serve as a foundation for future research on omni model reward modeling, evaluation, and post-training.
2 Related Work
Reward modeling for visual generation. Building on RLHF, reward modeling has been extended beyond text. ImageReward (Xu et al., 2023), HPSv2 and v3 (Wu et al., 2023b; Ma et al., 2025), PickScore (Kirstain et al., 2023) learn human preferences for text-to-image generation, improving correlation with human judgments and guiding diffusion models beyond CLIP-based proxies. For image editing, EditScore (Luo et al., 2025) and EditReward (Wu et al., 2025) adopt similar preference-learning frameworks. Unified approaches aim for cross-task generalization: Wang et al. (2025) train a single multimodal reward across image, video, and understanding tasks. Despite progress, most multimodal RMs remain task-specific and lack a unified, stress-testing evaluation.
Evaluating reward models. Benchmarking reward models has become an active research direction. In the text domain, RewardBench and RewardBench 2 (Lambert et al., 2025; Malik et al., 2025) systematically compare LLM reward functions across diverse axes (e.g., instruction following, reasoning, safety). VL-RewardBench (Li et al., 2025) and Multimodal RewardBench (Yasunaga et al., 2025) assess reward models for multimodal LLM. However, they only focus on image-text-to-text tasks. For image generation, researchers develope automatic evaluation metrics for text-to-image generation. CLIPScore (Hessel et al., 2021) offers a reference-free image–text similarity measure that correlates with human judgments but often misses compositional errors; TIFA (Hu et al., 2023), VQAScore (Lin et al., 2024) address this by probing alignment via VQA, improving robustness on compositional cases. The human annotations collected in these works are often used as reward model evaluations. Most existing reward model evaluations focus either on text or text-to-image generation, offering little insight into interleaved text and image. To bridge this gap, Multimodal RewardBench 2 (MMRB2) provides a unified and challenging framework for assess reward modeling for omni models.
3 Multimodal RewardBench 2
MMRB2 (Figure 3) is a comprehensive omni reward model evaluation benchmark spanning a range of tasks (§3.1) of four types: text-to-image generation, image editing, interleaved generation, and multimodal reasoning. Each datapoint in MMRB2 contains a task prompt (§3.2) and two model responses, chosen and rejected (§3.3). Reward models are evaluated based on their agreement with human annotators (§3.4).
3.1 Tasks in MMRB2
Figure 2: Breakdown of MMRB2 by task type and source, and detailed categories under each task.
Task 1. Text-to-Image. Text-to-image generation provides natural language prompts for which generators produce candidate images. Reward models see the prompt and the candidate images, and must prefer the human-preferred image based on factors such as object composition, spatial relationships, attribute binding, text rendering, and adherence to complex multi-object instructions.
Task 2. Image Editing. Image editing provides 1-3 input images and a textual edit instruction, along with candidate edited images from generators. Reward models must select the edit that best matches human preference, balancing faithfulness to the edit request with preservation of irrelevant regions. The edits include both single-image operations (e.g., changing attributes, scene modifications, adding/removing elements) and multi-image compositions where multiple inputs must be integrated.
Task 3. Interleaved Generation. Interleaved generation provides multimodal prompts that elicit mixed image–text sequences from generators (e.g., for storytelling, how-to guides, educational content, or multi-step reasoning). Reward models are asked to rank candidate interleaved outputs, capturing human preferences for coherence, global planning, and effective coordination between visual and textual content.
Task 4. Multimodal Reasoning (Thinking with images). Multimodal reasoning provides complex problems that require visual understanding, logical inference, and multi-step problem solving. Generators may produce both text and intermediate thinking or sketchpad images; reward models must judge which candidate reasoning trajectory and final answer better aligns with human preference, emphasizing accurate perception, spatial reasoning, and clear explanation.
See Figure 1 for examples of multimodal preference pairs in MMRB2 across these four subtasks.
3.2 Prompt and response collection
For each task, we sample prompts from existing benchmarks via stratified sampling over difficulty and subtask type, using only test splits to avoid train–test leakage. We additionally design new, practical tasks (e.g., multi-image editing) that are not covered in prior benchmarks. Benchmarks are weighted by coverage and difficulty, yielding 1,000 prompts per task. For each prompt, we generate multiple responses from 7–11 state-of-the-art models, including both API and open-source systems. We observe that even strong models such as Gemini 2.5 Flash Image struggle on interleaved generation and multimodal reasoning; for such cases, we further construct agents that can call Python and image generation/editing tools (Hu et al., 2024). Table 1 summarizes prompt sources and candidate models, with additional details in Appendix 9.
Figure 3: Overview of the MMRB2 data pipeline. The process combines ensemble MLLM judging, human verification, and multi-stage filtering to ensure high-quality, reasoning-consistent preference pairs across tasks.
3.3 Human preference annotations
Given prompts and responses, we developed methods to gather human preferences for each task type.
3.3.1 Image generation, editing & interleaved tasks
We adopt a unified annotation protocol to ensure consistency across text-to-image generation, image editing, and interleaved generation tasks.
Ensemble filtering. To focus human annotation on the most informative comparisons, we first apply an ensemble filtering pipeline that removes easy preference pairs where one response is almost unanimously preferred. We collect judgments from nine multimodal judges spanning API models (GPT-5, GPT-4.1, GPT-4o, Gemini 2.5 Flash, Gemini 2.5 Pro) and open-source VL models (Gemma-3-27B/12B/4B, Qwen-2.5-VL-7B), covering a wide range of capability. Each judge evaluates every pair twice, once in forward order (A vs. B) and once in reverse order (B vs. A), to mitigate position bias (see Appendix 9 for the exact prompts).
We define easy pairs as those where the majority label appears in at least 90% of all judge evaluations across both orderings, and discard them because they provide little signal about fine-grained differences between reward models. While ensemble filtering can in principle introduce bias, the diversity of the judges and the high 90% threshold restrict filtering to near-trivial cases and mitigate systematic bias from any single model.
Human Preference Annotation. We employed professional annotators via the Surge AI platform to collect high-quality human preferences. [^1] Annotators were compensated at an hourly rate of \$85. Each pair is independently evaluated by three annotators who have no knowledge of which model generated each response. Annotators assess each response using a comprehensive evaluation framework with different criteria tailored to each task category. Finally for each pair, annotators provide their overall preference for answer A vs B on a 7-point Likert scale where we convert these ratings to preferences. See details in Appendix 9. We implement several additional quality control measures. First, we filter out annotations with high inter-annotator disagreement, specifically removing pairs where the rating spread (maximum rating minus minimum rating) exceeds 4 points on the 7-point scale. We also exclude ambiguous annotations where the average rating falls too close to the scale midpoint (within the 3.0-4.0 range), as these indicate genuine uncertainty rather than clear preferences from human annotations. Finally, we remove pairs where the majority vote results in a tie, as these provide limited signal for evaluating judge agreement.
For the three generative tasks, we collected approximately 17,700 human preference judgments, each evaluated by three independent annotators, resulting in 5900 judgments overall. After filtering, we retain 1,000 pairs per task (approximately 50% of the initial set). Inter-annotator agreement on these filtered pairs is high: 95.6% overall (excluding ties), with task-specific rates of 95.3% for image generation, 96.3% for image editing, and 95.2% for interleaved generation.
3.3.2 Multimodal reasoning task
Because multimodal reasoning prompts have ground truth answers, we collect human annotations per model response (rather than pairwise) then construct pairs.
Human annotation. We filter generated model responses from §3.2 to those that contain both the correct answer to the prompt and some form of reasoning. We then balance samples across responses that include text-only reasoning and those that reason with both images and text. We collect three human annotations per response that indicate whether the reasoning contained in the model response is correct. The annotator instructions are listed in Appendix 8.2.
Pair construction. With the annotated responses, we construct preference pairs. For the human-preferred sample of each pair, we select model responses in which all three human annotators agree that the reasoning contains no major errors and the model answer is correct. For the non-preferred sample of each pair, we utilize two kinds of responses: Correct answer, incorrect reasoning, where the model answer is correct but all three annotators consider the reasoning to contain major errors, and Incorrect answer, with reasoning, where the model answer is incorrect and some form of reasoning is included. For each pair, the two model responses may share the same modality (both text-only or both image+text) or be a combination. No model response is duplicated across pairs. For more details, see Appendix 8.2.
3.4 Evaluation Method
Finally, we use the preference pairs to evaluate reward models on MMRB2.
Positional consistent dual evaluation Position bias is common problem; models have a systematic preference for the first item in a pair (Min et al., 2022; Tan et al., 2025). To mitigate this, each pair is evaluated twice per judge: once in its original order (A vs. B) and once with responses swapped (B vs. A). Both forward and reverse judgments are retained as independent data points, doubling judge-human comparison instances. This protocol improves agreement statistics by increasing sample size and penalizes judges with high position bias.
Judge-Human Agreement Computation We measure judge-human agreement by comparing each judge evaluation against the human preference for the corresponding pair. Human preference is determined by majority vote across three annotators for Tasks 1-3 and unanimous agreement of reasoning and answer correctness in Task 4. For each judge evaluation (whether forward or reverse), we compute a binary agreement score: 1.0 if the judge’s preference matches the human preference (including tie-to-tie matches), and 0.0 otherwise.
4 Experiments
We conduct a comprehensive study of omni reward modeling with MMRB2 along a number of dimensions: evaluation of MLLM-as-a-judge (§4.1), evaluation of other task-specific evaluators (§4.2), and in-depth analysis on various aspects of the benchmark and omni model reward modeling (§4.3 - 4.5).
4.1 Performance of MLLM-as-a-judge
Judge
Text to
Image
Image
Editing
Interleaved
Generation
Multimodal
Reasoning
Avg.
Open-source multimodal LLMs
Gemma 3 4B (Gemma Team et al., 2025) 51.7 51.0 51.3 48.8 50.7
Gemma 3 12B (Gemma Team et al., 2025) 56.0 58.0 58.0 49.3 55.3
Gemma 3 27B (Gemma Team et al., 2025) 58.3 60.2 61.1 49.4 57.3
Qwen2.5-VL-7B (Bai et al., 2025) 50.4 57.1 48.4 47.5 50.9
Qwen2.5-VL-72B (Bai et al., 2025) 59.1 64.6 62.3 50.0 59.0
Qwen3-VL-8B (Qwen Team, 2025) 59.4 61.7 61.5 54.6 59.3
Qwen3-VL-32B (Qwen Team, 2025) 64.1 67.3 70.5 56.6 64.6
Qwen3-VL-30BA3B (Qwen Team, 2025) 60.0 59.5 57.3 57.3 58.5
Qwen3-VL-235BA22B (Qwen Team, 2025) 62.0 64.8 69.0 55.9 62.9
API-based Models
GPT-4o (OpenAI, 2024) 60.3 65.0 61.5 51.9 59.7
GPT-4.1 (OpenAI, 2025a) 65.8 68.2 67.0 53.0 63.5
GPT-5 (OpenAI, 2025b) 70.5 73.8 74.4 70.2 72.2
Gemini 2.5 Flash (Gemini Team, 2025) 63.1 66.5 69.4 57.5 64.1
Gemini 2.5 Pro (Gemini Team, 2025) 70.5 71.3 75.1 66.6 70.9
Gemini 3 Pro (Google DeepMind, 2025b) 74.4 74.9 76.4 79.5 76.3
Table 2: MLLM-as-a-judge accuracies on MMRB2. The best numbers are bolded and the second best are underlined. Gemini 3 Pro is the best across all tasks. Qwen3-VL-32B is the best open-source model.
Setup. We evaluate all tasks on API-based models GPT-4o, GPT-4.1, GPT-5, Gemini 2.5 Flash, Gemini 2.5 Pro, Gemini 3 Pro and open-source models Qwen 2.5-VL (7B and 72B) (Bai et al., 2025), Qwen 3-VL (8B, 32B, 30BA3B, 235BA22B) (Qwen Team, 2025) and Gemma 3 (4B, 12B, and 27B) (Gemma Team et al., 2025). For each task type, we design task-specific evaluation prompts with detailed rubrics (see Appendix 9). We follow the positional consistent dual evaluation method in §3.4 to mitigate positional bias.
Results. Table 2 reveals substantial variation in judge-human agreement across models and tasks. API-based models generally outperform open-source alternatives, with Gemini 3 Pro achieving the strongest overall performance across all tasks. GPT-5 and Gemini 2.5 Pro also achieves decent accuracy on text-to-image generation, image editing, and interleaved generation (70 - 75% accuracy). Notably, multimodal reasoning proves to be the most challenging task across all models except Gemini 3 Pro, with even top API models achieving only 52-70% agreement on reasoning tasks (compared to 63-75% on multimodal generation tasks). This difficulty may stem from multiple valid solution paths, varying levels of explanation detail that humans may value differently, or the challenge of assessing both correctness and reasoning quality simultaneously.
Gemma 3, Qwen2.5-VL, and Qwen3-VL families of models all perform better on MMRB2 as number of parameters scales. Additionally, the performance gap between API-based and open-source models has narrowed with recent open-source advances. The top API models (Gemini 3 Pro, Gemini 2.5 Pro, GPT-5) achieve agreement rates of 65-80% across most tasks, while the best open-source models now reach competitive performance levels. Qwen3-VL-32B is the strongest open-source model, achieving 64.1-70.5% across tasks. Notably, its 70.5% agreement rate for interleaved generation approaches API-based model performance. While the Qwen3-VL series generally outperforms the Gemma 3 and Qwen2.5 families on image-related tasks, even some of the Gemma 3 and Qwen2.5 variants are within a few percentage points of API-based models. However, open-source models still show large gaps with API-based models on multimodal reasoning: the strongest, Qwen3-VL 30BA3B at 57.3%, trails Gemini 3 Pro by 22 percentage points.
4.2 Performance of supervised reward models
Judge
Text to
Image
Image
Editing*
Multimodal
Reasoning*
MLLM-as-a-judge
Qwen2.5-VL-7B (Bai et al., 2025) 50.4 57.8 53.7
Qwen3-VL-32B (Qwen Team, 2025) 64.1 66.4 69.9
GPT-5 (OpenAI, 2025b) 70.5 74.3 83.8
CLIP-based evaluators
CLIPScore (Hessel et al., 2021) 51.0 - -
ImageReward (Xu et al., 2023) 54.0 - -
HPSv2 (Wu et al., 2023a) 54.7 - -
PickScore (Kirstain et al., 2023) 58.6 - -
Qwen2.5-VL-7B-based evaluators
VQAScore (Lin et al., 2024) 58.3 - -
HPSv3 (Ma et al., 2025) 60.2 - -
EditReward (Wu et al., 2025) - 67.2 -
UnifiedReward (Wang et al., 2025) 59.8 - 55.1
Table 3: Other evaluators’ accuracies on MMRB2. Note that all task-specific evaluators except CLIPScore and VQAScore have been trained with human preference pairs. *For editing we use the single-image subset; for reasoning we use the text-only-output subset, ensuring fair comparison among evaluators.
Besides directly prompting MLLMs to act as judges, prior work has proposed a range of automatic metrics and preference-trained reward models targeting the tasks in MMRB2. We evaluate these methods on the three MMRB2 tasks—text-to-image generation, image editing, and multimodal reasoning. To the best of our knowledge, there are currently no evaluators specifically tailored for interleaved text–image outputs.
Setup. Unless otherwise noted, we adopt the default metaprompt provided by each official library. For text-to-image, we consider two families of evaluators. The first is CLIP-based (Radford et al., 2021), including CLIPScore (Hessel et al., 2021) and its preference-trained variants ImageReward (Xu et al., 2023), HPSv2 (Wu et al., 2023a), and PickScore (Kirstain et al., 2023). The second family is based on Qwen2.5-VL-7B (Bai et al., 2025). We evaluate VQAScore (Lin et al., 2024), which scores generated images using model logits, as well as the preference-trained reward models HPSv3 (Ma et al., 2025) and UnifiedReward (Wang et al., 2025). We evaluate all of the above models on the MMRB2 text-to-image task. Qwen2.5-VL-7B has also been used as the backbone for reward models on other tasks, including EditReward (Wu et al., 2025) for image editing and UnifiedReward (Wang et al., 2025) for multimodal understanding. Because EditReward is trained only on single-image editing, and UnifiedReward is trained on single-image image-to-text tasks, we evaluate them on the corresponding single-image subsets of MMRB2 to ensure a fair comparison among evaluators. Table 3 summarizes the results.
Preference training substantially improves reward-model accuracy. Several reward models share the same base architecture as our MLLM baselines (e.g., EditReward, UnifiedReward, and HPSv3 are based on Qwen2.5-VL-7B), and some are CLIP-based (ImageReward, HPSv2, PickScore). Relative to the Qwen2.5-VL-7B judge, EditReward yields a +9.4% gain on editing (57.8 $\rightarrow$ 67.2), and UnifiedReward improves text-to-image by +9.4% (50.4 $\rightarrow$ 59.8) and reasoning by +1.4% (53.7 $\rightarrow$ 55.1). Similarly, compared to CLIPScore (51.0), CLIP-based preference models show consistent gains: ImageReward 54.0 (+3.0 %), HPSv2 54.7 (+3.7 %), and PickScore 58.6 (+7.6 %). These results indicate that training with human preferences is an effective way to boost evaluator performance on multimodal tasks.
Reward models can be out-of-distribution; strong MLLMs remain strong judges. Despite the above gains, most preference-trained reward models still underperform a larger open-source judge such as Qwen3-VL-32B across tasks; a notable exception is EditReward, which is competitive on editing (67.2 vs. 66.4). One plausible explanation is a distribution shift: several reward models were trained on data from earlier-generation systems (e.g., SD 2.1–era), and their accuracy diminishes when judging outputs from more capable, recent models. Overall, newer reward models (HPSv3, EditReward, UnifiedReward) are far better than older ones, yet stronger MLLM still set a high bar through simple prompting.
4.3 Correlation with downstream tasks
Figure 4: Downstream best-of-N score v.s. MMRB2 performance. We perform best-of-N sampling with 2 base models each on 4 tasks (GenAI-Bench (Li et al., 2024), GEdit-Bench (Liu et al., 2025), ISG-Bench (Chen et al., 2024), and EMMA (Hao et al., 2025)). A judge’s score on MMRB2 strongly correlates with improvement in downstream tasks when it is used in best-of-N sampling, highlighting MMRB2’s utility for downstream task success.
A key research question is whether MMRB2 performance can predict downstream task performance. To address this question, following prior works (Lightman et al., 2023; Li et al., 2025), we conduct best-of-N sampling with different rewards. We experiment with 4 tasks: GenAI-Bench (Li et al., 2024), GEdit-Bench (Liu et al., 2025), ISG-Bench (Chen et al., 2024), and EMMA (Hao et al., 2025), each corresponds to one task in MMRB2. For each query, we generate $N=8$ candidate responses from two base models, one strong model and a weaker one, and we use 7 different MLLM-as-a-judge to select the best one via knockout tournaments. Then we evaluate the selected response with each task’s metrics.
Figure 4 shows all the results. The x-axis is the MMRB2 performance, and Y-axis is the score of the best-of-N response selected by different rewards. We exclude GPT-4o and 4.1 because they are often used as evaluators. For interleaved generation, we remove ISGBench preference pairs when computing MMRB2 scores to avoid leakage. The results show that there is a strong correlation between best-of-N performance and MMRB2 scores. A good reward model can give great gains on downstream performance, even with the simple best-of-N sampling. For example, FLUX’s GenAI-Bench score improved from 73% to 79%, and GPT-4o’s accuracy on EMMA improved from 32% to 45%, when using GPT-5 as best-of-N selector. We can still see consistent gains even for strong base models like Gemini 2.5 Flash Image and o3-Sketchpad. The strong correlation (>0.8 Pearson’s $r$ for all tasks and models) between MMRB2 and downstream task performance validates that MMRB2 is a good proxy for downstream effectiveness.
4.4 Fine-grained analysis of errors
Same-model pairs vs. different-model pairs. Our benchmark contains 57.4% same-model pairs (comparing two outputs from the same model) and 42.6% different-model pairs (comparing outputs from different models), allowing us to assess judge performance across both scenarios. See results in Table 4.
Task Judge Overall (%) Same-M (%) Diff-M (%)
Image Generation Gemini 3 Pro 74.4 70.4 79.7
Gemini 2.5 Pro 70.5 68.4 73.4
GPT-5 70.5 66.8 75.6
GPT-4.1 65.8 61.6 71.4
Qwen3-VL-32B 64.1 59.1 71.0
Image Editing Gemini 3 Pro 74.9 71.0 79.3
GPT-5 73.8 71.7 76.2
Gemini 2.5 Pro 71.3 66.7 76.6
GPT-4.1 68.2 65.6 71.3
Qwen3-VL-32B 67.3 64.5 70.5
Interleaved Gemini 3 Pro 76.4 72.8 82.0
Gemini 2.5 Pro 75.1 70.7 81.9
GPT-5 74.4 69.4 82.1
Qwen3-VL-32B 70.5 66.7 76.4
Gemini 2.5 Flash 69.4 65.0 76.3
Reasoning Gemini 3 Pro 79.5 78.7 79.8
GPT-5 70.2 68.4 70.8
Gemini 2.5 Pro 66.6 70.5 65.4
Gemini 2.5 Flash 57.5 59.9 56.7
Qwen3-VL-30BA3B 57.3 57.4 57.3
Table 4: Detailed performance breakdown of top 5 judges per task showing overall agreement, same-model pairs, and different-model pairs with human preferences.
Across the image generation, editing, and interleaving tasks, we observe a consistent pattern for all judges: judges achieve higher agreement with human on different-model pairs compared to same-model pairs. For the best-performing judges, this gap ranges from 5-13 percentage points. For example, on image generation, Gemini 3 Pro achieves 79.7% agreement on different-model pairs but only 70.4% on same-model pairs (9.3 point gap). This pattern holds across tasks: same-model pairs demand fine-grained judgments within one model’s outputs, while different-model pairs reveal larger gaps rooted in capability differences.
Same-modality pairs vs mixed-modality pairs. For the multimodal reasoning task, we study how judges perform differentially for pairs constructed with responses from the same modality (e.g., text response vs. text response) versus mixed modalities (e.g., text response vs. text-image response). Full results are reported in Table 5.
Same modality:
Image+text
Same modality:
Text
Mixed modality:
Pref: Image+text; Not Pref: Text
Mixed modality:
Pref: Text; Not Pref: Image+text
Model
Correct vs.
incorrect reason
Correct vs.
incorrect answer
Correct vs.
incorrect reason
Correct vs.
incorrect answer
Correct vs.
incorrect reason
Correct vs.
incorrect answer
Correct vs.
incorrect reason
Correct vs.
incorrect answer
Open-source models
Gemma3 4B 47.39 50.63 50.00 48.32 63.64 57.00 38.68 36.00
Gemma3 12B 49.57 51.47 54.02 52.10 81.82 73.50 15.09 11.50
Gemma3 27B 51.30 50.21 51.79 51.68 87.50 79.50 10.38 10.50
Qwen2.5-VL-7B 49.12 48.10 51.34 50.00 52.27 39.00 47.17 40.31
Qwen2.5-VL-72B 52.63 48.10 53.57 54.41 78.41 68.00 16.04 23.98
Qwen3-VL-8B 57.46 52.53 58.48 54.20 71.59 73.00 34.91 36.73
Qwen3-VL-32B 62.28 54.43 60.71 56.93 78.41 80.00 25.47 33.16
Qwen3-VL-30BA3B 58.77 55.72 57.59 56.30 75.00 78.00 36.79 43.37
Qwen3-VL-235BA22B 58.11 57.02 55.80 57.14 85.23 81.96 23.58 26.02
API-based models
GPT-4o 50.43 50.42 55.80 56.51 81.82 80.00 18.87 18.00
GPT-4.1 56.09 50.42 58.04 58.61 93.18 81.50 10.38 13.50
GPT-5 69.57 67.02 75.89 80.25 88.64 88.00 36.79 40.00
Gemini 2.5 Flash 60.53 58.47 56.25 59.03 86.36 76.00 16.98 38.42
Gemini 2.5 Pro 73.91 66.18 62.95 65.55 84.09 79.00 43.40 58.00
Gemini 3 Pro 71.88 84.75 75.45 82.49 84.88 87.00 66.98 72.00
Table 5: Multimodal reasoning performance breakdown by pair modality and pair type.
We find that for mixed-modal pairs, all judges exhibit a strong bias towards the response that contains images. This is true even of the highest performing models: the accuracy of GPT-5 for mixed-modal pairs when the preferred response contains an image is 49.3 points higher than pairs where the preferred response contains text (88.2% vs. 38.9%), and Gemini 2.5 Pro and Qwen3-VL-30BA3B have gaps of 27.7 and 36.0 points respectively. Gemini 3 Pro performs much better, but still has a 17.9 point gap. Additionally, we find that this trend holds for both pair types: those constructed with an incorrect response vs. a correct response and those with incorrect reasoning vs. correct reasoning.
4.5 Test-time scaling of rewards
Figure 5: Majority-vote accuracy of each MLLM as the number of samples $K$ varies. Test-time scaling yields small gains for GPT and Gemini models but no improvement for Qwen3-VL.
Prior work (Wang et al., 2022a; Brown et al., 2024) shows that test-time scaling can substantially improve LLM performance. We ask whether similar gains transfer to multimodal reward models. For each judge, we draw $K\in\{1,3,5,7,9\}$ independent judgments and take a majority vote as the final decision. We report majority-vote accuracy averaged over the four MMRB2 tasks (300 examples per task) in Fig. 5.
[…]
5 Conclusion
We introduce Multimodal RewardBench 2, the first comprehensive benchmark for omni reward models spanning four tasks: text-to-image, image editing, interleaved generation, and multimodal reasoning. Our analysis suggests that current omni reward models, particularly the latest Gemini 3 Pro, can serve as proxies for human evaluation on multimodal generation tasks, achieving 74-80% agreement. However, the substantial disagreement remaining (20-26%) indicates that human evaluation remains essential, , and that other models, including GPT-5, lag significantly behind Gemini 3 Pro. Overall, MMRB2 establishes a challenging and informative benchmark that we hope will serve as a foundation for future research on omni model reward modeling, evaluation, and post-training.
Limitations and future extensions. As the first comprehensive benchmark targeting omni reward models, MMRB2 focuses on core settings and overall human preferences in interleaved text–image scenarios. The construction pipeline is modular and can be extended to additional evaluation dimensions (e.g., safety- and bias-sensitive preferences), richer task formats (e.g., multilingual tasks, in-the-wild prompts, multi-turn and agentic interactions), and further modalities (e.g., video and audio). Further discussion is provided in Appendix 10.
6 Acknowledgements
We would like to thank Mason Yu, Christophe Ropers, Nate Ekberg, Cynthia Gao, Justin Hovey, Jaimie Hsu, Samantha Snowden, and all annotators from Surge AI for their invaluable contributions to data annotation. We also thank Jonea Gordon and Vanessa Stark for their assistance with the approval process. Additionally, we are grateful to Mary Williamson, Xiaochuang Han, Adriana Romero Soriano, Michal Drozdzal, Xudong Wang, Michihiro Yasunaga, Ishan Misra, Amita Kamath, Inna Lin, and Karen Chen for their insightful discussions and support throughout this project.